RC meeting and Run Plan, Monday January 23rd
Return to the RC Meeting Calendar
Contents
RC Meeting Information
Current RC: | Iuliia Skorodumina | |
Time: | Daily at 3:30PM | |
Location: | Counting House Conference Room |
Remote Connection Details
- Join ZoomGov Meeting
- https://jlab-org.zoomgov.com/j/1619032184?pwd=aWhTVWxya1ZZQzA1K2ZPSTd2ZX...
- Meeting ID: 161 903 2184
- Passcode: 260715
- Dial by your location
- +1 669 254 5252 US (San Jose)
- +1 646 828 7666 US (New York)
- +1 551 285 1373 US
- +1 669 216 1590 US (San Jose)
- 833 568 8864 US Toll-free
- Find your local number: https://jlab-org.zoomgov.com/u/acj3pTicMx
Summary
Sunday Swing (01/22). Shift started with 30 uA, then we went up to 35 uA.
At aroung 17:30 there was a power glitch (https://logbooks.jlab.org/entry/4115123). The glitch affected SBS magnet, which tripped. Before recovering from the first power glitch, another occurred (https://logbooks.jlab.org/entry/4115152), so there was no beam for about 3 hours.
When the SBS magnet tripped, no alarm followed (https://logbooks.jlab.org/entry/4115142). The Shift Crew discovered the magnet issue when they observed abnormalities during the NMR Measurement (https://logbooks.jlab.org/entry/4115137). The issue with the SBS magnet trip affected the target polarization, which dropped a bit (down to 42.6% at 18:00). The polarization quickly recovered, though, rising up to 48.2% by 20:30.
Once the accelerator recovered from the power glitches, we performed a quick beamcheck (https://logbooks.jlab.org/entry/4115180) and around 20:30 went back to production. MCC started with delivering 35 uA, then we went up to 40 uA, and at around 21:30 up to 45 uA.
There were several GEM HV trips after the beam restoration. GEM expert was contacted and the issue was fixed (https://logbooks.jlab.org/entry/4115195 and https://logbooks.jlab.org/entry/4115224).
Swing Shift summary (https://logbooks.jlab.org/entry/4115103).
Monday Owl (01/23). We were taking production data at stable 45 uA overnight. However, not without difficulties.
BBCal HV issue (https://logbooks.jlab.org/entry/4115349). It turned out that the HV values for problematic channels were not loaded, so Arun helped the Shift Crew to load them (https://logbooks.jlab.org/entry/4115366).
CODA crashed and was fixed by Alex (https://logbooks.jlab.org/entry/4115340).
GEMs HV tripped and Sean had to increase the trip threshold (https://logbooks.jlab.org/entry/4115384).
Finaly, at the end of the shift the target IOC crashed again and Jack had to reboot it (https://logbooks.jlab.org/entry/4115472).
Owl Shift summary (https://logbooks.jlab.org/entry/4115391).
Monday Day (01/23). The shift started with one more CODA crash around 9am, which took about an hour to fix. Once the CODA was fixed, the beam went down for about 3 hours. At about 13:30 we did a quick beam checkout (https://logbooks.jlab.org/entry/4115697) and went back to production at 45 uA.
Jiwan pointed out that HCal TDC errors reappeared this morning. He rebooted HCAL roc17 crate and the errors seem to be gone (https://logbooks.jlab.org/entry/4115747). Jiwan requests the Shift Crew to check on TDC errors. See also Jiwan's older logentry on how to notice the TDC issues (https://logbooks.jlab.org/entry/4110023). See also an example of TDC warnings in 50k replay (https://logbooks.jlab.org/entry/4110346).
Day Shift summary (https://logbooks.jlab.org/entry/4115780).
We plan on going to Controlled Access tomorrow morning (Tuesday Day, 01/24).
Run plan
- Continue production on pol He3 cell at 45 uA.
- Tuesday Day (01/24) -> MCC starts beam studies at 8:30 am (for 4 hours) -> Controlled Access?
- Moller Measurement? January 30 is a new scheduled date for the Wien angle change because Hall B will not start running until then.
Other issues
- CODA keeps crashing.
- Should we add magnets alarms into the Alarm Handler?
Shift Schedule
Please consider signing up for shifts [1].